Please do everything neetly

5. An open tank of very large cross section area cross section area a A is filled with water to a height 6.5 cm2 tank allows bott n the bottom of the tank allows water to dran id at is the volume flow rate at which water flows out b) At what distance below the bottom of the tank is the cross- one-half of the area of the hole? If A-10000a, how long will it take to drain the tankthe If there were two holes with equal cross section areas flow rate and the time of drain change? c) d)
Answer #1

velocity of water coming out=(2gh)^.5

=2*10*.3^.5

=2.45m/s

Volume flow rate=area*velocity

=6.5*245

=1592.1cm^3/s

time to empty=A/a(2h/g)^.5

=10000(2*.3/10)^.5

=2450s

If 2 holes are made volume flow rate will double and the time to empty tank will become half of the value found above